OpenStudy (anonymous):

Please help i am really stuck ! Find three consecutive integers such that the sum of the first and twice the third is one more than three times the second. Thanks for all help!!! :)

OpenStudy (anonymous):

what are consecutive integers? :3

OpenStudy (cwrw238):

let the 3 integers be n , n+1 and n+2 then we have n + 2(n + 2) = 3(n + 1) + 1 solve for n

OpenStudy (cwrw238):

but this is an identity !!???

OpenStudy (cwrw238):

so there are infinite solution

OpenStudy (cwrw238):

are you sure you have the question right/

OpenStudy (anonymous):

yea idk maybe it was a misprint

OpenStudy (cwrw238):

as it stands you can fit any number to n 1,2,3.....

OpenStudy (cwrw238):

- or negative ones for that matter

OpenStudy (texaschic101):

n = 1st integer n + 1 = 2nd integer n + 2 = 3rd integer n + 2(n + 2) = 3(n + 1) + 1 n + 2n + 4 = 3n + 3 + 1 3n + 4 = 3n + 4 I agree...infinite solutions